A 19 g sample of an alloy at 98.0°C is placed into 84.6 g of water at 22.0 °C in an insulated coffee cup with a heat capacity of 9.2 J/K. If the final temperature of the system is 35.0°C, what is the specific heat capacity of the alloy in J/(g.K)? Don't include units. cH2O = 4.184 J/g.K
Answer #1

m(water) = 84.6 g

T(water) = 22.0 oC

C(water) = 4.184 J/goC

m(alloy) = 19.0 g

T(alloy) = 98.0 oC

C(alloy) = to be calculated

C3 = 9.2 J/oC

We will be using heat conservation equation

use:

heat lost by alloy = heat gained by water and 3

m(alloy)*C(alloy)*(T(alloy)-T) = m(water)*C(water)*(T-T(water)) + C3*(T-T(water))

19.0*C(alloy)*(98.0-35.0) = 84.6*4.184*(35.0-22.0) + 9.2*(35.0-22.0)

1197*C(alloy) = 4601.5632 + 119.6

C(alloy)= 3.94 J/goC

C(alloy)= 3.94 J/gK

Answer: 3.94 J/gK